-5y+6=-34
I need you're help with this for me PLZ can u help me !I have my question in the top PLZ HELp me !

Answer by checkley77(12844)

-5y+6=-34
-5Y=-34-6
-5Y=-40
Y=-40/-5
Y=8 ANS.